OpenStudy (anonymous):

MEDAL AND FAN Justify if completing the square is a good method for solving when the Discriminant is negative. Use any of your three functions as an example and respond in complete sentences. f(x) = x^2 - 4 g(x) = x^2 + 6x + 3 h(x) = x^2 + 2

OpenStudy (anonymous):

I have a similar question, I do not know if you can complete the square with only two terms?

OpenStudy (anonymous):

You can in at least some cases. But if there's a better method then it would be a waste of time to attempt completing the square. Using f(x) as an example, I notice that's also a difference of squares, so I wouldn't complete the square there.

OpenStudy (anonymous):

Can I use that as my answer?

OpenStudy (anonymous):

For f(x), sure, but you'd have to continue to explain why it's more efficient. Your reasoning for h(x) and g(x) can't be the same, though.

OpenStudy (anonymous):

Where do negative discriminants come into play? I don't get that.

OpenStudy (anonymous):

h(x) is actually supposed to be h(x) = x^2 + 3x + 6

OpenStudy (anonymous):

I'd complete the square for that one.

OpenStudy (zzr0ck3r):

the square is "completed" already on two terms, complete the square means get it into the form of (x+a)^2+b, and you have this already with x^2+2 x^2+2 = (x+0)^2+2 so its already in vertex form....no need to do anything, just set equal to 0 and solve, if you are trying to find solutions.

OpenStudy (anonymous):

I don't know what my answer should be :/ lol

OpenStudy (anonymous):

@mathmale Hey could you help me?

OpenStudy (mathmale):

I'd be glad to. Which problem would you like to focus on?

OpenStudy (anonymous):

@mathmale Sorry, I fell asleep. Could we try h(x) = x^2 + 3x + 6?